OpenStudy (anonymous):

A Norman window has the shape of a semicircle atop a rectangle so that the diameter of the semicircle is equal to the width of the rectangle. What is the area of the largest possibe Norman window with a perimeter of 36 feet?

OpenStudy (anonymous):

hope this helps !! Area of rectangle = (2r)*h = 2rh Area of semicircle = 1/2 pi r^2 So we want to maximize A = 2rh + 1/2 pi r^2 but there are two variables, so we'll first try to eliminate one of them. Note that perimeter = 2h + 2r + pi r = 31 so h = 31 - r - pi/2 r = 31 - (1 + pi/2)r Substitute that into our equation for area to get: A = 2r(31-(1+pi/2)r) + 1/2 pi r^2 = 62r - 2r^2 - 1/2 pi r^2 Take the derivative: dA/dr = 62 - 4r - pi r and set it equal to zero: 0 = 62 - 4r - pi r (4 + pi) r = 62 r = 62 / (4 + pi) Now that you have the radius, you can find the area A by using the equation A = 62r - 2r^2 - 1/2 pi r^2

OpenStudy (anonymous):